OpenStudy (anonymous):

Solve the equation. |k + 6| = 3 A. {–3, 9} B. {–3, 3} C. {–9, –3} D. {all real numbers greater than or equal to –9 and less than or equal to –3} I think it's C but can someone show me how to solve this?

OpenStudy (anonymous):

@ilikemath50

OpenStudy (anonymous):

You are correct

OpenStudy (anonymous):

thanks but how is solved?

OpenStudy (anonymous):

|k+6|=3 is: k+6=3 and k+6=-3. solve for k in both to get {-9,-3}

OpenStudy (anonymous):

awesome thanks mate =D

OpenStudy (anonymous):

You welcome bruh!